Tickets for an event cost $4 for children, $12 for adults, and $7 for senior citizens. The total ticket sales were $1920. There were 50 more adult tickets sold than child tickets, and the number of senior citizens tickets were 4 times the number of child tickets. How many of each ticket were sold?
Answer by stanbon(48554)

Tickets for an event cost $4 for children, $12 for adults, and $7 for senior citizens. The total ticket sales were $1920. There were 50 more adult tickets sold than child tickets, and the number of senior citizens tickets were 4 times the number of child tickets. How many of each ticket were sold?
-------------
Let number of child tickets be "x"; their value is 4x dollars.
Number of adult tickets is "x+50" ; their value is 12(x+50)=12x+600 dollars
Number of senior tickets is "4x" ; their value is 7(4x)=28x dollars
--------------
EQUATION:
value + value + value = 1920 dollars
4x+12x+600+28x=1920
44x=1320
x=30 (# of child tickets)
x+50=80 (# of adult tickets)
4x=120 (# of senior tickets)
Cheers,
Stan H.